# Modeling when and how physics PhD students search for a research group: the role of interests and prior research experiences in timely group integration

### v2

@@ III.3 Data analysis
Once completed, interviews were transcribed and edited for grammar and clarity. The transcripts became the subject of our thematic analysis, which followed the steps outlined in Braun and Clarke [81]. Analysis began with repeated and active reading of the data and application of a priori codes aligned with the stages of the protocol (e.g., “Step” or “Question”). For each, we also applied an initial code describing the excerpt and what we found important about it, which is a common first step in grounded theory approaches to data analysis. Initial coding is an open-ended first cycle coding process that involves breaking down qualitative data into discrete parts, examining them closely, and applying codes that promote deep reflection on the contents of the data [82]. Our initial codes typically consisted of a short phrase or sentence describing the excerpt and what we found important about it.
@@
At this point we developed a codebook specifically focused on the research alignment aspect of students’ narratives. We coded excerpts along two broad axes: comments regarding students’ personal Interest development, and statements about how students characterized the process of Evaluating research alignment. Excerpts coded under Interest development were indicative of how well a student knew what research they wanted to pursue in graduate school. The code Evaluating research alignment was applied to comments in which students described what characteristics they were looking for in a research group with regard to congruence of research interests. In some cases, both codes were applied to an excerpt. For instance, the excerpt “I was looking at programs to apply for… and if I couldn’t find people that did theoretical stellar astrophysics I crossed them off the list” was double coded to capture the interest shown in theoretical astrophysics and the activity of evaluating programs’ available research.
As the coding process continued, these codes were refined and broken into new themes and sub-themes. The results of our analysis are presented in Section IV. In particular, results stemming from our analysis of student Interest development are presented in Section IV.2 and results from our analysis of student efforts when Evaluating research alignment are presented in IV.4.
@@
In addition to thematically organizing the data, we also linked student-level meta data information to all 40 transcripts regarding when each student committed to joining their current research group. While many students had a clear idea of when they committed to their research group (e.g., “I shook on it with my PI, saying I was going to commit to joining the lab fully.”), this was not always the case. For instance, some students described having an interest in a group before graduate school and beginning to perform research activities with the group in their first semester, but never described a distinct moment when they joined the group. Approximately $N=7$ students presented this ambiguity. The authors (M.V.) and (B.Z.) discussed these cases individually. Together they agreed on the time when these students most strongly began to characterize themselves as members of their group, typically evidenced by participation in group research activities and statements about their confidence in which research topic they wanted to pursue. For example, one student never described a precise moment when she joined her research group, but said that after she met her current advisor during her senior summer, “I don’t think I ever really thought to look for a different advisor.” However, we classified this student as joining her group in her first semester of graduate school when she actually began to perform research activities in this advisor’s group. Our analysis of when the physics PhD students in our sample joined their research groups is discussed in Section IV.1.
@@
We used several strategies to produce a valid and reliable coding scheme. One researcher (M.V.) was responsible for developing most of the codes and themes, which were refined through discussion between (M.V.) and (B.Z.) at weekly research meetings. Inter-rater reliability was done once the final codebook was established in order to demonstrate its validity. In particular, a selection of 40 random excerpts corresponding to students’ extent of interest development and stage of evaluating research were given to a researcher unaffiliated with the project to code using the sub-themes discussed in the results. Agreement was high, with disagreements primarily occurring over borderline cases (e.g., whether a student who had primarily looked at only a few research groups before arriving to graduate school but had never contacted them would count as evaluating school level research alignment or evaluating individual group level alignment). We refined the codebook after discussion to more clearly delineate when certain codes should be applied.
III.4 Analysis of program handbooks
@@
Lastly, we sought to determine each program’s requirements regarding when student should (or must) join a group, or whether programs had any requirements about this subject at all. To do so, we visited each program’s website to find resources about their PhD timelines and program requirements. These were commonly located in a graduate student handbook. Only 2 of the 13 schools did not have an online graduate student handbook available. One school did not appear to have any publicly available resources for expected PhD timelines and program requirements. For analysis, we examined the resources to find the first mention of students needing an advisor. We collected these instances and grouped them based on whether finding an advisor was an explicit or implicit requirement, as well as the specificity of when in students’ PhD timeline those requirements occurred. Results are presented in Section IV.1.
@@ IV.1 When physics PhD students join research groups and the perceived benefits of joining sooner
Time of interview
When committed to group 1st year fall 1st year spring 2nd year fall 2nd year spring Total
Senior year 1 0 0 0 1
Senior summer 5 1 0 1 7
1st semester 4 3 1 2 10
[] Subtotal: 1st semester or before 10 4 1 3 18
2nd semester 1 2 2 4 9
1st year summer 0 1 3 3 7
2nd year or later 0 0 2 1 3
Has not joined 2 1 0 0 3
[] Subtotal: 2nd semester or after 3 4 7 8 22
Total 13 8 8 11 40
Table 2: The semester that students in our sample reported committing to their graduate research group, broken down by which semester of graduate school they were interviewed. $N=18$ students joined in their first semester or before, which implies that much of their search took place before arriving to graduate school. $N=22$ joined in their second semester or later, aligning with more typical timelines suggested in graduate handbooks. Students who joined a group later tended to have been interviewed later as well.
@@
Lovitts found that PhD students who establish the adviser-advisee relationship early in their graduate careers are more likely to complete their PhDs successfully. To investigate this phenomenon in the context of physics graduate programs, we start by looking at when students in our sample joined their groups. Using an analysis of program handbooks, combined with Lovitts’ observation that the second semester is a critical period during which students begin to consider leaving, we motivate our decision to categorize students who joined a group in their first semester or before as early joiners. We then illustrate some of the perceived benefits experienced by those students who found a group quickly, contrasting them with narratives of students who committed to a research group later on.
@@
Programmatic expectations: Analysis of departmental expectations of when students should join a group reveals strongly varying levels of guidance across the programs studied. Only one provided a specific time by which PhD students were required to have found an advisor (the fall semester after passing a written qualifying examination). On the other hand, 5 schools structured their expected PhD timelines around completion of courses and qualifying exams, and did not explicitly provide a time by which students were expected to join a group. Rather, joining a research group was implicitly needed in order to complete the other program requirements. For instance, one school required that students pass an oral exam on a topic jointly chosen by the student and their research advisor by the beginning of the third year. Finding an advisor is clearly necessary for this step, but it was not an explicit expectation in the timeline.
@@
Most commonly, programs provided a “typical timeline” or an “expected timeframe” for finding a research group. Six schools fell into this category, but their expectations varied. For example, one strongly suggested students should join a group in the spring semester of their first year, while another suggested that it was more typical to “formally” join a research group in the third year of the program. Another provided a broad swath of time (spring semester of first year to summer of the third year) during which finding an advisor was expected, but noted that earlier was better. The lone program that employed a formal laboratory rotation requirement for students expected students to have chosen their research advisor by the end of their first academic year. Despite ambiguity across most of the program handbooks regarding when students are expected to find an advisor, they generally indicate that a “typical” timeline for students to find an advisor should be around the end of their first academic year or beginning of their second year.
@@
When students join research groups: The total number of students in our sample who joined a group during or after their 2nd semester of graduate school was $N=22$. 16 of the 22 students joined within the expected period of most PhD programs, while three joined in their second year or later. Three students had not joined a group at the time of their interviews, but these students were either already in their second semester of graduate school or were nearly finished with their first semester. Thus, we categorized their timelines with those students who joined after their second semester.
Notably, the time to join a group reported in Table IV.1 represents the time that students took to join their most current research group. However, several had joined earlier than the time indicated in the table and subsequently chose to switch. $N=7$ of the 22 students who reported finding a group later in their graduate programs had done so after switching out of a different group. For a majority of students, switching research groups was a personal decision associated with a unique set of challenges. Thus, we chose not to include the three students in our sample whose programs had formal rotations. Although these students changed research groups during their first year, it was a mandatory part of their program rather than a decision they made on their own.
@@
In order to have joined a group in their first semester of graduate school or before, those students likely conducted much of their search before they arrived in their programs. This is particularly true of the students who started working with their advisors the summer before their programs started, who were well ahead of the expected timelines published by their universities. On the other hand, students who joined in their second semester or later more likely searched for groups during their first year of graduate school, and their timelines align more closely with what program handbooks describe as more “typical.”
@@
Perceived benefits of joining early: Students who joined a research group before or during their first semester discussed several perceived benefits. Several students noted that getting into a group before the first year of coursework reduced anxiety about their future due to the security it offered them. Nina expressed how finding a group “did take a lot of weight off my shoulders… not necessarily having to spend too much time obsessing about oh, everybody else already has a group.” Luis had two offers for research groups before his first year of graduate school began, which he described as an advantage over other students in his class: “I felt so relieved that whole first year, that I didn’t have to go out and look for a group as actively as everyone else did right?… a lot of people really struggled to find a good fit, whereas I felt like, ‘Oh, I could send an email right now and get it in the group that I want, or in a group that wants me.”’ In both of these quotes, students describe comparing their own progress to other students in their cohort, which eased their anxiety. However, the reverse was true for students who did not join a group right away. Dev, who struggled to determine if his preferred advisor would be able to fund him or not, said during his second semester of graduate school “It is a bit stressful though in the sense that, for some of the people, before even coming here they had fixed on an advisor. Everything was fixed… they’ll go and do the PhD. So talking to [those students] was not really helpful because they are not going through any of this.”
@@
Others noted that getting into research faster made them feel like they were making progress toward their degree. Heather credited joining a group shortly after entering graduate school for her accelerated graduation timeline: “if I had not been in [my advisor’s] group as a first year and kind of set on that path, I don’t think I would have had the resources or knowledge to be ready to [take my qualifying exam] that that early.” Being able to partake in research during her first year allowed Heather to take her research-focused qualifying exam earlier than most other students in her class. This then allowed her to apply for several grants that were only available to students who had completed their qualifying exam, furthering her research agenda. On the other hand, Jack had not joined a group at the time he was interviewed in his second semester. He was hoping to join a group that he perceived as highly competitive, and he had taken the professor’s elective class during his first semester to demonstrate his interest. Jack was concerned about being behind his desired PhD timeline if he was unable to join his preferred group at the end of his second semester: “I’ve put in a lot of time into this, and if it ends up not working out, it’s a bit annoying because of, yeah, just wasted time. I’d say it’s a bit stressful for sure.”
@@
The comments in this section demonstrate the effect that the time students join a group can have on their emotional state and overall outlook on how well they are progressing through their PhD program. Moreover, these themes broadly align with our prior findings that students who struggled to find a research group tended to feel a lower sense of belonging in their programs, and suggests the importance of continuing to study the factors that influence when students are able to join a research group. In subsequent sections, we characterize several factors that lead some students to more quickly navigate the group search process in physics. In particular, we reveal several major differences between these students regarding level of interest development, when they started evaluating individual research groups, and the kinds of difficulties they experienced during their group search process.
@@ IV.2 Some physics PhD students know what research they want to do, but many do not
Figure 2: We conceptualize physics PhD students’ certainty regarding their research interests as existing on a spectrum. For this analysis, we grouped students into three categories: broad interests, interest in subfield and method, and knowing exact research. Students with broad interests were less sure of what research they wanted to pursue in graduate school, while students who knew the exact research they wanted often already had a research group in mind upon entering graduate school.
@@
Since students’ experiences searching for a research group often spanned several semesters, some interviewees described how their interests shifted over time. Some students made statements coded as Broad interests during their junior year of undergrad, but later described how other factors (e.g., research experiences, talking with advisors) caused them to become interested in particular topics. These students would have excerpts appear in multiple categories of interest, but at different times in their narrative. Thus, the sum of students who made comments coded under each of the categories is higher than the total number of students in the sample. Figure 3 breaks down each category to elucidate how many students were uniquely coded under a particular sub-code, or whether they appeared under multiple categories. For example, the figure illustrates that 15 students only had excerpts coded as Broad interests, indicating that they had entered graduate school open to a variety of research topics. Meanwhile, a different 6 had excerpts in both the Broad interest and Interest in subfield and method categories, indicating that these students became more confident in their interests over time.
@@
Other students described more specific indecision, such as whether they wanted to pursue more or less hands-on work. For example, Hassan was interested in topics related to quantum computing, but noted that his small undergraduate institution did not offer enough research opportunities for him to determine the kind of research he wanted to do. He said, “I didn’t really have a lot of research background. So like, in undergrad, I got involved with like one project, which was small scale given the school that I attended. But I was not really sure what I wanted to do, theory or experimental research.” Cole similarly had interests in quantum-related topics, but did not have a strong preference for the kind of research he wanted. He described searching for graduate schools based on the idea that “I wanted to do things with quantum stuff, maybe use machine learning or something, do something wacky.” However, as Cole later noted, many subfields of physics involve “quantum stuff” and use computational techniques such as machine learning, leaving numerous research pathways open to him.
A significant subset of these students ($N=8$) were still trying to narrow down their interests during or after their first semester in the graduate program. Kali recalled that during her second semester of graduate school she still “wasn’t sure exactly what I wanted to do… again, I came in not knowing anything.” Meanwhile, Greg remarked that by his second semester in graduate school he had not yet decided whether he wanted to pursue biophysics or condensed matter: “I was open to doing whatever research in grad school, whether it be biophysics or condensed matter. I wasn’t like, heart set on one.” Isaac noted that by his second semester in graduate school, “I didn’t really know what I was interested in… all I knew was, like, something that involved quantum to some extent.” Thus, these students entered graduate school without a clear idea of which research they wanted to pursue and had not clarified their decision during their first semester. More often, students were too busy to worry about finding a group. Per Isaac, “First semester, finding an advisor was kind of like the last thing on my mind… [friends] would ask me sometimes, have you found a research group? And I’d be like no, I haven’t even thought about it. I’m trying to figure out the electric field of a charged hemisphere, or whatever the homework was about.”
Figure 3: Diagram illustrating the number of unique students who had excerpts coded under each interest code. Since students’ narratives spanned multiple semesters, they often expressed how their interests developed over the course of their timeline. For example, the 6 in the diagram indicates that six students had described having broad interests at one time, but later had become more interested in a particular subfield and method. Notably, 15 students only ever discussed having broad interests. Numbers represent unique students, so their sum is 40.
@@
Knowing exact research A smaller subset of graduate students in our sample ($N=13$) reported being highly confident in the research that they wanted to pursue, and often knew the exact research group they wanted to join. For instance, Heather said that being able to work on machine learning and supernovae was “the one thing that was actually keeping me in physics and astronomy” (Senior year). Indeed, upon entering graduate school and beginning to work in her current research group, she affirmed, “This is exactly what I want to be doing in astronomy. Like, if I decide this specific thing isn’t for me, there’s not really anything else I’d rather be doing here, y’know?” Hence, she suspects that if things “didn’t work out” with her current advisor, “I probably wouldn’t stay in grad school. Like, at that point, I would just go into industry.” Heather’s comments highlight how strongly some students’ research interests can influence their willingness to persevere in graduate school.
Most commonly, students described a strong interest in working with a particular faculty member at their prospective graduate school. For instance, Rose described already having a “person of interest” with whom she wanted to work in graduate school, while Lakshmi said she “had it pretty sorted in my mind where I wanted to go” before arriving. Selena recalled, “I’m pretty sure my whole admissions essay was just, I want to work with [this professor]. And also, these other people are cool, but mostly [him]. Yeah, I sort of came in very much angling for that.” She noted that this certainty was based primarily on this professor’s research topic, saying “I didn’t really know much about him other than just what he did. You know, like what’s available on a website for a professor. But it definitely was like, this is stuff I think is really cool.”

### v3

@@ III.3 Data analysis
Once completed, interviews were transcribed and edited for grammar and clarity. The transcripts became the subject of our thematic analysis, which followed the steps outlined in Braun and Clarke [87]. Analysis began with repeated and active reading of the data and application of a priori codes aligned with the stages of the protocol (e.g., “Step” or “Question”). For each, we also applied an initial code describing the excerpt and what we found important about it, which is a common first step in grounded theory approaches to data analysis [88]. Initial coding is an open-ended first cycle coding process that involves breaking down qualitative data into discrete parts, examining them closely, and applying codes that promote deep reflection on the contents of the data [89]. Our initial codes typically consisted of a short phrase or sentence describing the excerpt and what we found important about it.
@@
At this point we developed a codebook specifically focused on the research alignment aspect of students’ narratives. We coded excerpts along two broad axes: comments regarding students’ personal Interest development, and statements about how students characterized the process of Evaluating research alignment. Excerpts coded under Interest development were indicative of how well a student knew what research they wanted to pursue in graduate school. The code Evaluating research alignment was applied to comments in which students described the characteristics they were looking for in a research group with regard to congruence of research interests. In some cases, both codes were applied to an excerpt. For instance, the excerpt “I was looking at programs to apply for… and if I couldn’t find people that did theoretical stellar astrophysics I crossed them off the list” was double coded to capture the interest shown in theoretical astrophysics and the activity of evaluating programs’ available research. As the coding process continued, these codes were refined and broken into new themes and sub-themes. The results of our analysis are presented in Section IV. In particular, results stemming from our analysis of student Interest development are presented in Section IV.2 and results from our analysis of student efforts when Evaluating research alignment are presented in IV.4.
@@
In addition to thematically organizing the data, we also linked student-level meta data information to all 40 transcripts regarding when each student committed to joining their current research group. While many students had a clear idea of when they committed to their research group (e.g., “I shook on it with my PI, saying I was going to commit to joining the lab fully.”), this was not always the case. For instance, some students described having an interest in a group before graduate school and beginning to perform research activities with the group in their first semester, but never described a distinct moment when they joined the group. Approximately $N=7$ students presented this ambiguity. The authors (M.V.) and (B.Z.) discussed these cases individually. Together they agreed on the time when these students most strongly began to characterize themselves as members of their group, typically evidenced by participation in group research activities and statements about their confidence in which research topic they wanted to pursue. For example, one student said that after she met her current advisor during her senior summer, “I don’t think I ever really thought to look for a different advisor.” Still, we classified this student as joining her group in her first semester of graduate school when she actually began to perform research activities in this advisor’s group. Our analysis of when the physics PhD students in our sample joined their research groups is discussed in Section IV.1.
@@
We used several strategies to produce a valid and reliable coding scheme. One researcher (M.V.) was responsible for developing most of the codes and themes, which were refined through discussion between (M.V.) and (B.Z.) at weekly research meetings. Inter-rater reliability was done once the final codebook was established in order to demonstrate its validity. In particular, a selection of 40 random excerpts corresponding to students’ extent of interest development and stage of evaluating research were given to a researcher unaffiliated with the project to code using the sub-themes discussed in the results. Agreement was high, with disagreements primarily occurring over five borderline cases (e.g., whether a student who had primarily looked at only a few research groups before arriving to graduate school, but had never contacted them, would count as evaluating school level research alignment or evaluating individual group level alignment). We refined the codebook after discussion better delineate when certain codes should be applied.
Figure 2: (a) Summary of the main topics discussed in the results section. Although the impact of students’ background experiences on the group search is mentioned throughout the results, it is the focus of subsections IV.3.1 and IV.4.1. (b) Condensed version of the model presented in Section V, previewing the relationships between the variables that we establish throughout the results.
@@
Lastly, to gain more information about the institutional requirements regarding when students should (or must) join a group, we analyzed the graduate program handbooks of the schools represented in the sample. Only 2 of the 13 schools did not have an online graduate student handbook available. One school did not appear to have any publicly available resources for expected PhD timelines and program requirements. For analysis, we examined the resources to find the first mention of students needing an advisor. We collected these instances and grouped them based on whether finding an advisor was an explicit or implicit requirement, as well as the specificity of when in students’ PhD timeline those requirements occurred. Results are presented in Section IV.1. An overview of the results section and which major topics are covered in each section is given in Figure 2.
@@ IV.1 When physics PhD students join research groups and the perceived benefits of joining sooner
Lovitts found that PhD students who establish the advisor-advisee relationship early in their graduate careers are more likely to complete their PhD successfully [12]. However, the definition of “early” is unclear in the context of physics graduate education, as there is no prior literature documenting when students join research groups. By examining when students in our sample found their research groups, we motivate our decision to categorize students who joined a group in their first semester or before as “early joiners.” These students perceived several benefits from joining a group quickly, and noted that their experiences during the first year contrasted with those of students who committed to a research group later on. We also motivate our definition of “joining early” with an analysis of program handbooks to better understand when departments expect students to find a group.
@@
Programmatic expectations Analysis of departmental expectations of when students should join a group reveals strongly varying levels of guidance across the programs studied. Only one provided a specific time by which PhD students were required to have found an advisor (the fall semester after passing a written qualifying examination). On the other hand, 5 schools structured their expected PhD timelines around completion of courses and qualifying exams, and did not explicitly provide a time by which students were expected to join a group. Rather, joining a research group was implicitly needed in order to complete the other program requirements. For instance, one school required that students pass an oral exam on a topic jointly chosen by the student and their research advisor by the beginning of the third year. Finding an advisor is clearly necessary for this step, but it was not an explicit expectation in the timeline.
@@
Most commonly, programs provided a “typical timeline” or an “expected timeframe” for finding a research group. Six schools fell into this category, but their expectations varied. For example, one strongly suggested students should join a group in the spring semester of their first year, while another suggested that it was more typical to “formally” join a research group in the third year of the program. Another provided a broad swath of time (spring semester of first year to summer of the third year) during which finding an advisor was expected, but noted that earlier was better. The lone program that employed a formal laboratory rotation requirement for students expected them to have chosen their research advisor by the end of their first academic year. Despite ambiguity across most of the program handbooks regarding when students are expected to find an advisor, they generally indicate that a “typical” timeline for students to find an advisor should be around the end of their first academic year or beginning of their second year.
When committed to group N
Senior year 1
Senior summer 7
1st semester 10
[] Subtotal: 1st semester or before 18
2nd semester 9
1st year summer 7
2nd year or later 3
Has not joined 3
[] Subtotal: 2nd semester or after 22
Total 40
@@
Table 2: The semester that students in our sample reported committing to their graduate research group. $N=18$ students joined in their first semester or before, which implies that much of their search took place before arriving to graduate school. $N=22$ joined in their second semester or later, aligning with more typical timelines suggested in graduate handbooks. Three students had not joined a group at the time of their interviews, but these students were either already in their second semester of graduate school or were nearly finished with their first semester. Thus, we categorized their timelines with those students who joined after their second semester.
When students join research groups The total number of students in our sample who joined a group during or after their 2nd semester of graduate school was $N=22$. 16 of the 22 students joined before their second year began, which falls within the expected timeframe of most PhD programs. Meanwhile, three joined in the middle of their second year or later. Three students had not joined a group at the time of their interviews, but these students were either already in their second semester of graduate school or were nearly finished with their first semester. Thus, we categorized their timelines with those students who joined after their second semester.
@@
In order to have joined a group in their first semester of graduate school or before, those students likely conducted much of their search before they arrived. This is particularly true of the students who started working with their advisors the summer before their programs started, who were well ahead of the expected timelines published by their universities. On the other hand, students who joined in their second semester or later more likely searched for groups during their first year of graduate school, and their timelines align more closely with what program handbooks describe as more “typical.”
Notably, the time to join a group reported in Table IV.1 represents the time that students took to join their most current research group. However, several had joined earlier than the time indicated in the table and subsequently chose to switch. $N=7$ of the 22 students who reported finding a group later in their graduate programs had done so after switching out of a different group. For a majority of students, switching research groups was a personal decision associated with a unique set of challenges. Thus, we chose not to categorize the three students in our sample whose programs had formal rotations as having “switched.” Although these students changed research groups during their first year, it was a mandatory part of their program rather than a decision they made on their own.
@@
Students distinguish joining earlier or later Students who joined a research group before or during their first semester discussed several perceived benefits compared to their peers. Several students noted that getting into a group before the first year of coursework reduced anxiety about their future due to the security it offered them. Nina expressed how finding a group “did take a lot of weight off my shoulders… not necessarily having to spend too much time obsessing about oh, everybody else already has a group.” Luis had two offers for research groups before his first year of graduate school began, which he described as an advantage over other students in his class: “I felt so relieved that whole first year, that I didn’t have to go out and look for a group as actively as everyone else did right?… a lot of people really struggled to find a good fit, whereas I felt like, ‘Oh, I could send an email right now and get it in the group that I want, or in a group that wants me.”’ In both of these quotes, students describe comparing their own progress to other students in their cohort, which eased their anxiety. However, the reverse was true for students who did not join a group right away. Dev, who struggled to determine if his preferred advisor would be able to fund him or not, said during his second semester of graduate school “It is a bit stressful though in the sense that, for some of the people, before even coming here they had fixed on an advisor. Everything was fixed… they’ll go and do the PhD. So talking to [those students] was not really helpful because they are not going through any of this.” Dev’s quote highlights feeling like an outsider compared to other students who seemed to be further ahead.
@@
Others noted that getting into research faster made them feel like they were making progress toward their degree. Heather credited joining a group during her first semester of graduate school for her accelerated graduation timeline: “if I had not been in [my advisor’s] group as a first year and kind of set on that path, I don’t think I would have had the resources or knowledge to be ready to [take my qualifying exam] that that early.” Being able to partake in research during her first year allowed Heather to take her research-focused qualifying exam earlier than most other students in her class. This then allowed her to apply for several grants that were only available to students who had completed their qualifying exam, furthering her research agenda. On the other hand, Jack had not joined a group at the time he was interviewed in his second semester. He was hoping to join a group that he perceived as highly competitive, and he had taken the professor’s elective class during his first semester to demonstrate his interest. Jack was concerned about being behind his desired PhD timeline if he was unable to join his preferred group at the end of his second semester: “I’ve put in a lot of time into this, and if it ends up not working out, it’s a bit annoying because of, yeah, just wasted time. I’d say it’s a bit stressful for sure.”
@@
The comments in this section motivate our choice to demarcate the first semester or before as an “early” time to join a group. They also begin to demonstrate the effect that the time students join a group can have on their emotional state and overall outlook on how well they are progressing through their PhD program. Those themes broadly align with our prior findings that students who struggled to find a research group tended to feel a lower sense of belonging in their programs [32]. In subsequent sections, we characterize several factors that lead some students to more quickly navigate the group search process in physics. In particular, we reveal several major differences between students regarding level of interest development, when they started evaluating individual research groups, and the kinds of difficulties they experienced during their group search process.
Figure 3: A summary of codes used in Results Section IV.2. We conceptualize physics PhD students’ certainty regarding their research interests as existing on a spectrum. For this analysis, we grouped students into three categories: broad interests, interest in subfield and method, and knowing exact research. Students with broad interests were less sure of what research they wanted to pursue in graduate school, while students who knew the exact research they wanted often already had a research group in mind upon entering graduate school.
@@ IV.2 Some physics PhD students know what research they want to do, but many do not
Since students’ experiences searching for a research group often spanned several semesters, some interviewees described how their interests shifted over time. Some students made statements coded as Broad interests during their junior year of undergrad, but later described how other factors (e.g., research experiences, talking with advisors) caused them to become interested in particular topics. These students have excerpts appear in multiple categories of interest, but at different times in their narrative. Thus, the sum of students who made comments coded under each of the categories is higher than the total number of students in the sample.
@@
Other students described more specific indecision, such as whether they wanted to pursue more or less hands-on work. Hassan recalled that coming into graduate school, “I was kind of not really sure what I wanted to do, theory or experimental research.” Ash had nearly the same sentiment, recalling that “Coming in, I was deciding if I wanted to do high energy theory or high energy experiment.” Meanwhile, Cole described having interests in quantum-related topics, but did not have a strong preference for the kind of research he wanted. He described searching for graduate schools based on the idea that “I wanted to do things with quantum stuff, maybe use machine learning or something, do something wacky.” However, as Cole later noted, many subfields of physics involve “quantum stuff” and use computational techniques such as machine learning, which left numerous research pathways open to him.
A significant subset of students who reported broad interests ($N=8$) did so during or after their first semester in the graduate program. Kali recalled that during her second semester of graduate school she still “wasn’t sure exactly what I wanted to do… again, I came in not knowing anything.” Greg remarked that by his second semester, he was still “open to doing whatever research in grad school, whether it be biophysics or condensed matter. I wasn’t like, heart set on one.” Meanwhile, Isaac noted that in his second semester, “I didn’t really know what I was interested in… all I knew was, like, something that involved quantum to some extent.” Thus, these students entered graduate school without a clear idea of which research they wanted to pursue and had not clarified their decision during their first semester. Most often, these students were too busy with other responsibilities to worry about finding a group. Per Isaac, “First semester, finding an advisor was kind of like the last thing on my mind… [friends] would ask me sometimes, have you found a research group? And I’d be like no, I haven’t even thought about it. I’m trying to figure out the electric field of a charged hemisphere, or whatever the homework was about.”
@@
Knowing exact research A smaller subset of graduate students in our sample ($N=13$) reported being highly confident in the research that they wanted to pursue. For instance, Heather said that being able to work on machine learning and supernovae was “the one thing that was actually keeping me in physics and astronomy.” Indeed, upon entering graduate school and beginning to work in her current research group, she affirmed, “This is exactly what I want to be doing in astronomy. Like, if I decide this specific thing isn’t for me, there’s not really anything else I’d rather be doing here, y’know?” Hence, she suspects that if things “didn’t work out” with her current advisor, “I probably wouldn’t stay in grad school. Like, at that point, I would just go into industry.” Heather’s comments highlight how strongly some students’ research interests can influence their willingness to persevere in graduate school.
Most commonly, students described a strong interest in working with a particular faculty member at their prospective graduate school. For instance, Rose described already having a “person of interest” with whom she wanted to work in graduate school, while Lakshmi said she “had it pretty sorted in my mind where I wanted to go” and who she wanted to work with before arriving. Similarly, Selena recalled that “I’m pretty sure my whole admissions essay was just, I want to work with [this professor]. And also, these other people are cool, but mostly [him]. Yeah, I sort of came in very much angling for that.” She noted that this certainty was based primarily on this professor’s research topic, saying “I didn’t really know much about him other than just what he did. You know, like what’s available on a website for a professor. But it definitely was like, this is stuff I think is really cool.”
While Selena’s interest in this faculty member was driven primarily by her perceived interest alignment, others attributed their confidence in wanting to work with a particular faculty to other reasons as well. Ursula had worked with a particular faculty member through a collaboration in her undergraduate research, which gave her confidence in her ability to work in his group. “The head of [the collaboration’s] gravitational waves group is actually the person that I applied to work with here.” She continued, explaining how her prior experience informed her that he would be a good mentor: “I knew that if I were to work with him, I’d have to kind of relearn a lot of things… but it felt like with his guidance, and with his wealth of knowledge, and willingness to explain things, and excitement about explaining things, it felt really worth it.”